OMC Closing Review (2022/11/18)
Higher mode rejection -> OK with good mode matching. -> will be checked after pumping.
- New optical board to remove additional structure for resonance dumping.
New high-quality mirror. -> 97% T_FP.
- Replacement of the broken PD
800 Hz resonance still remains. -> unstable lock in air. Pumping down OMC are will mitigate this resonance.
- FP_tra estimation in air ~ 0.89 derived by subtracting HOMs power sum from the total power.
- OMC alignment. OSTM should not be moved by pico to avoid trouble.
- Dark noise: same btw both PDs.
OFI performance check to investigate mysterious power loss (60% !!). -> Good enough! No problem.
PD protector: Fast beam shutter(we don't have) or PZT offset to unlock the OMC. -> Urgent issue. 50mW/one PD is acceptable.
- Update plan to improve PD circuit for the in-air lock?
Notch filter cannot help to avoid resonance -> 800Hz is in the obs band!! Another circuit for the in-air work?
Why DC-PD was broken? -> It was damaged during removing the glass window. A high-quality can opener is necessary. A can opener of Thorlabs that has a shaper edge.
- Quantum Efficiency of PD ~ 9? %. No actual measurement. Uncertainty is large. Measure the QE of a spare PD?
- The calibration factor is included in MEDM.
- Transfer Function
- OSTM has a different TF(LPY) compared with OMMT1/2 because it has a thinner TM.
OSEM output values vary with each other. Coupling force variation. -> LPY coupling will be generated. Maybe acceptable because we don't use them for control.
- Space btw magnets and plates
- Took photos. All gaps are OK.
- Pico motion
TF of H1~4 phases are OK? Actuation signs (directions) are managed well? -> OSTM has different of magnets S-N direction after magnet repair.
Positive -> push? Should be confirmed.
Push -> OSEM voltage should be negative.
- Transfer Function
- OMMT2T optical setting for sensing OMMT1/SR2 motion separately (Ushiba, Akutsu, Miyoki, Sato, Takahashi, Yokozawa)
- Vacuum Issues
Clamp preparation -> OK, f300 -> metal, f800 -> elastomer,
TMP/IP alignment OMC-OMMT-TMIPunit.pdf
- Opened viewports confirmation
- Cleaning inside
- Leak in GV_srm-OMM again??
- Loosen the tierod bolts to free the GV motion to the OMM side.
- Close GV btw BS-IXYC for the emergency.
- Retighten bolts of the support of the duct near the GV_IYA.
Slow pumping as IFI-IMMT/IMC -> Suspension Monitoring(Takahashi-san + alpha)